1.     Bacteria A and Bacteria B are considered hardy bacteria and can survive in extreme environments when grown individually. However, when they are grown together in a closed environment (on an agar plate), Bacteria A thrives, while Bacteria B eventually dies off. 

a)     Calculate the density per cubic centimetre of Bacteria A on the agar plate if there are 1500 bacteria on a 100cm3 plate. (1 mark)

 

Dp=N/A

1500/100^3

= 15?


is this right??????


1
Expert's answer
2021-04-01T06:57:36-0400

The density of bacteria equals:

d = n / V

where d - density, n - number of bacteria, V - agar volume

From here:

d= 1500/1003 = 15 cm-3

So, the answer is correct.